Curriculum Vitae - ROBERT ANDREW DAVIS

Date of birth:  25th May 1970  Address: Churchill College,
Nationality:  British  Cambridge, CB3 0DS.
Marital status:  Single  Tel:  0223-323406 (home)

0223-332614 (work)

EDUCATION

1991–  University of Cambridge (Churchill College)

PhD in Control Engineering; “Model Validation for Robust Control” . During my PhD I have reviewed papers submitted for publication in various journals and supervised undergraduates.

1988–1991 University of Warwick

BSc (Hons.) in Mathematics, class I.

1986–1988 Bilborough College, Nottingham

A-levels: Maths (A^{1}), Further Maths (A),Physics (A), Chemistry (A).

1981–1986 Glaisdale School, Nottingham

Ten O-Levels (including Maths and English): 9 grade A, 1 grade B.

WORK EXPERIENCE

Summer 1991 Division and Waterfront Head at Summer Camp in USA. I was responsible to the Camp Director for the forty two children and twelve staff members in my division. As the Waterfront Head I was responsible for all the watersport activities, and the swimming teaching, of my division.

Summer 1990 Cabin Counsellor in USA. I was responsible for the eight boys in my cabin

during their time at camp and the coaching of swimming and athletics.

Summer 1989 Teaching maths in a Comprehensive School.

Casual work in postal department.

Summer 1988 Casual work in silk screen printers.

ACTIVITIES & ACHIEVEMENTS

Cambridge University • SERC sponsored course introducing project management, team- working and negotiation skills.

• Shell vacation course giving insight into manufacturing engineering and management.

• Organised activities on a residential camp for handicapped adults and underprivileged children.

• Secretary of the University swimming club.

• Half Blue for swimming.

• Rowed and played waterpolo for college.

Curriculum Vitae - ROBERT ANDREW DAVIS

Warwick University • Captain of the University swimming team.

• University Colours for swimming.

• Represented the University in swimming, athletics and biathlon.

• Wrote sports reports for the University newspaper.

School  • Pupil representative on school governing body.

• City schools 400m athletics champion 1986-88.

CURRENT INTERESTS

Coaching Having qualified as a swimming coach in 1993 I now enjoy coaching the University

team. I also coach children and adults with the City of Cambridge Club.

Sports I swim competitively for both the University and City of Cambridge. I also enjoy

rowing, squash and golf.

Travel Working in America gave me the opportunity to travel around most of the country during my three summers there. Since then I have travelled around northern Portugal.

OTHER QUALIFICATIONS/SKILLS

Computing Knowledge of UNIX, BASIC, PASCAL, MATLAB, ASSEMBLER and experi-

ence of using various micro-computers and work stations.

Other  Full clean driving licence.

REFEREES

Professor Keith Glover (PhD Supervisor)  Robert Grabill (Camp Director)
Cambridge University Engineering Department Camp Pemigewassett
Trumpington Street  Wentworth
Cambridge  NH 03282
CB2 1PZ  USA